This is my homework problem: "$500 was originally invested into an account that earns simple interest. after 3 years the account had a total worth of $650. Write an equation to model the situation and then Find the interest rate."
I have a partial equation of I = (500)(r)(3)
Or 650 + 500 = R
My biggest issues is that I honestly do not understand at all what I'm doing with the percent and interest equations. I don't know which one is correct.
I Do know that over the three years $150 came into the account at a rate of $50 a year...the problem is I don't know how to convert that knowledge into a percentage.
Thank you so much for solving this problem for me.

Answer by ankor@dixie-net.com(22740)

"$500 was originally invested into an account that earns simple interest.
after 3 years the account had a total worth of $650.
Write an equation to model the situation and then Find the interest rate."
:
The interest formula I = P*r*t
In this problem:
I = 150
P = 500
r = interest rate (which we are to find)
t = 3 yrs
so we have:
150 = 500 * r * 3
150 = 1500r
r = 150/1500
r = .10 which is 10% interest
:
You can confirm this in the interest equation
I = 500 * .10 * 3
I = 150
:
Did this make things clear in your mind? You have to agree once you see it, it is not very hard.